OpenStudy (anonymous):

Ms amy is a park ranger. She wants to estimate the number of fifh in a small lake. she captures and marks 50 fish. she returns them to lake . several week later she captures another 50 fish.she notices that 15 fish of the fish are some of the ones she marked. write o proportion that ms amy could use to predict the nnumber of fish in the lake? estimate number of fish in the lake?

OpenStudy (anonymous):

m=15/50 = 3/10 n=(3/10 ) * 50 = approx. 166.67 which is approx 167 fish